# Generalization of Task Parameterized Dynamical Systems using Gaussian Process Transportation

### v1

@@ V-B Multiple Reference Frames
Fig. 9 shows the box plot that compares the performance of the different models, i.e., TP-GMM, HMM, DMP, and GPT, on the training set. Nine demonstrations are available for different configurations of the starting and goal frame. When training the GMM models, i.e., TP-GMM and HMM, a subset of demonstration $m$ is randomly chosen from the training set and compared with the remaining $(9-m)$ demonstrations when evaluating the model in that unknown situation; the number of used demonstration is highlighted as an apex, e.g., HMM_6 means that we used an HMM model with six demonstrations. When training the transportation models, linear or non-linear, i.e., DMP or GPT, only one demonstration is randomly picked from the training set and compared with the other eight unseen situations. For each model, the random selection of demonstration and comparison is repeated $20$ times. Five metrics are used to compare the rollout trajectory and the actual demonstration:
@@
Considering that we have many models that can behave differently according to the amount and quality of the demonstration, it is not straightforward to deduce any conclusions on which method is statistically better from the boxplot of Fig. 9. For this reason, we run a U test, also known as the Mann-Whitney non-parametric test [38], to deduce if the distribution of results of each of the methods is statistically lower ($p<0.05$) than each of the others. When computing the U test between two methods, in case of a statistical difference, the winning method gets one point. The numbers on top of the figure for each of the methods indicate the performance ranking, i.e., the method that obtained the most points when computing the U-test is going to be first in the ranking. When more methods share the same position in the ranking, it simply means they were significantly better than the same amount of other methods during the comparisons.
@@
Figure 8: Qualitative comparison of multi-reference frame parameterization. Comparison between HMM, TP-GMM, DMPs, and our proposed method. The first raw compared the performance in the reproducing of the training set demonstration, depicted as a dashed black line, where both HMM and TP-GMM are training using all the nine demonstrations while DMP and GPT are only trained with one demonstration (the central one) and generalized for each of the frames. In the second row, a random perturbation is applied to each frame, and the model is queried on the most likely trajectory. For GPT, the uncertainty in the generalization is depicted with the orange areas. The blue stars are the points tracked during the motion, given that our proposed method does not rely on reference frames but only on source and target points.
Figure 9: Box plot results and performance ranking on frame configuration from the training set. The number on top of each box plot is the position of the method in the performance ranking.
V-C Multi-source single-target generalization
Fig. 8, in the column of Gaussian Process Transportation, depicts the generalization of a single demonstration from one single 2-frame source to multiple 2-frame targets. Although, in Fig. 1, we already depicted the generalization of many demonstrations and the learned dynamics from one source surface to another target, we still did not mention the generalization from multi-source to a single target. When dealing with n source distributions, we fit n different transportation functions $\phi$, each trained with different source points but the same target points.
Fig. 8 highlights how the many demonstrations given in different frame configurations can be transported in the same target frame and how we can extract a reactive policy, encoded in the vector field, as a function of the global position [25].
Figure 11: Multi-source single target generalization. Demonstrations from different frame positions (see Fig. 8 are transported on a single target multi-frame configuration ( unknown from the training set). The dashed line is the given human demonstration in that configuration. The vector field is the resulting dynamics learned also with a Gaussian Process with minimization of uncertainties from [25].
VI Real Robot Validation
To validate the proposed method on real manipulation tasks, we selected three challenging tasks, i.e., robot reshelving (Sec. 12), dressing (Sec. VI-B) and cleaning (Sec. 16), to teach as a single demonstration and generalize it in different scenarios. These are all tasks where the training set will never be similar to the test set; for example, when dressing a human, the configuration and shape of the arm may change, and we expect the robot to generalize the behavior accordingly.
We controlled a Franka Robot using a Cartesian impedance control [^3] https://github.com/franzesegiovanni/franka_human_friendly_controllers.
The dynamics of the demonstrations are learned with a non-parametric function approximation for motor learning that uses a joint position-time encoding, proposed in [39]. The distance from the next attractor position and orientation are a function of the current robot position. Our goal is to show how the proposed transportation theory can correctly generalize the pose, velocity, and stiffness of the robot. The following sections will summarize the robot validation experiments. A video of all the experiments can be found at https://youtu.be/FDmWF7K15KU.
VI-A Robot Reshelving
@@
Robot reshelving refers to picking an object in one location, moving it, and placing it in a desired position on a shelf.
Our assumptions for the problem are:
  • one global frame dynamical system is learned from a single demonstration and transported in the different object/goal configurations;
  • corner points of the objects and the shelf slot are tracked rather than position/orientation.
Fig. 12 depicts the experimental setup where a milk box, with an AprilTag [40] on it, has to be positioned on a compartment on a shelf, also marked by another tag. Before the demonstrations or execution, the robot searches for any frames in the spaces using the camera attached to its end-effector. For every frame, the transportation policy extracts a cube’s center and corners with predefined side dimensions as the markers. Fig. 12 shows how the demonstration for reshelving on the left of the central compartment can be generalized to any other floor, both on the left and right. We randomized the object position and orientation and the goal on the shelf ten different times, all successfully generalized. Table II shows the range of x,y,z, and yaw angles of the object and goal markers during the ten different executions, while Fig. 13 depicts the relative position w.r.t. the object and the frame of the different rollouts; from the figure it is possible to appreciate how the execution lines converge on the (initial) object position when picking and on the goal position when placing the object.
@@
Figure 13: Relative position of the end-effector w.r.t. the initial object and the goal position during multiple generalizations rollouts in robot reshelving.
Figure 14: Dressing policy generalization. Cyan marker is the shoulder, magenta is the elbow, yellow is the wrist, and blue is the hand. The blue rollout end-effector trajectory starts from the red dot and finishes with the green dot. $\alpha$ is the angle of the elbow; the smaller the angles, the more complicated the generalization would be.
VI-B Robot Dressing
The task of dressing is a primary task in elderly care. It consists of pulling a deformable sleeve over the posture of a human arm.
@@ VI-B Robot Dressing
VI-C Robot Surface Cleaning
@@
One of the main advantages of the proposed method is that it does not need to reconstruct the surface but only learns the map from the source to the target pointcloud. The deformation between the source and the target surface point cloud is modeled using a Stochastic Variational Gaussian Process Transportation (SV-GPT) to generalize the demonstrated policy position, orientation, and stiffness profile for a successful cleaning task. Given the large number of points in the source and target point cloud, i.e., 400 points, using a reduced set of inducing points, i.e., 100, makes fitting the transportation model more computationally efficient.
@@
Fig. 15 depicts the teaching of a cleaning task on a flat surface and the generalization on different higher, titled, and curved surfaces that belong to common objects. The lower row shows what the robot perceives of the environment; the blue dots in space are the source distribution, recorded before giving the demonstration (depicted as dashed line), and target distributions recorded before executing the roll-out transported policy (depicted a solid line). Fig. 15 also highlights how the roll-outs follow the shape of the surface, showing a successful generalization of the robot position and orientation. As previously stated, no external force-torque sensor is used to adapt the orientation of the end-effector on the tangential direction of the surface. However, an observer of the applied external force between the robot and the surface is estimated from measured torques in the joints. Fig. 16, depicts the estimated norm of the force exchanged with the surface, where the same increasing/decreasing trend is captured on the different surfaces.
VII Limitations and Open Challenges
@@
For example, we assume knowing the matching between the points of the source and the target distribution. However, in many complex scenarios, this limitation can be problematic, and some different pre-processing algorithms, such as optimal transport[41] or iterative closest point (ICP), need to be adopted to perform the matching. Additionally, semantic matching can increase cross-domain generalization, for example, by adapting the reshelving strategy to a completely different shelf type or adjusting the dressing policy from an adult to a baby arm.
@@
Another assumption of the developed method is dealing with static environments, i.e., the target distribution does not change during the policy’s rollout. However, this assumption can fall when dealing with the reshelving of moving objects or when trying to dress real humans that will probably move before and during the interaction. Nevertheless, supposing to know the state of the target distribution, the transportation policy can be updated inexpensively online since only the label $\hat{y}$ of Eq. (7) and the (cheap) linear transformation needs to be recomputed. However, the fitting of the transported policy $\hat{f}$ makes it challenging to perform the generalization online.
@@
Finally, given that in complex scenarios, the generalization may be inaccurate, the use of interactive human corrections may increase the resulting manipulation performance [1]. However, changing the generalized policy opens the question of whether interactive corrections should be propagated back to the source policy and how. Additionally, in case many source distributions/policies are recorded, the choice of generating the target policy by transporting all of them, like in Fig. 8, or by selecting the best one, according to some similarity criteria, can open exciting developments of the proposed theory.
VIII Conclusions
@@
In this paper, we address the prominent but challenging problem of policy generalization to novel unseen task scenarios. We formulate a novel policy transportation theory that, given a set of matched source and target points in the task space of the robot, regresses the function that, most likely, would match the source and target distribution. Additionally, we showed how the same transportation function and its derivatives can be exploited to transport the original policy dynamics, rotation, and stiffness while retaining uncertainties in the process. The same algorithm, which uses a Gaussian Process at its core, was tested and compared with different state-of-the-art regressors or different generalization methods, showing how, even with only one source demonstration, it results in better or comparable performance. However, the main requirement, for a successful generalization is to track and match important task points in the original scenario, where the demonstration was given, and the corresponding points in the new scenarios.
@@
We validated the proposed approach on a Franka Robot, testing it on three different tasks: product reshelving, arm dressing, and surface cleaning. These various tasks were never tackled together by the same generalization algorithm, and they usually were performed with ad hoc solutions, for example, to keep a constant force when cleaning a surface. Despite this, our policy transportation algorithm performed successfully in all of them. The tracking requirements were satisfied using fiducial markers or directly the point cloud estimated with the infrared camera sensor. Future development will have to focus on scaling the process on big and unmatched point clouds of complex (and deformable) objects to manipulate while allowing the use of human feedback in the fine-tuning of the resulting policy.

### v2

@@ III-H Transporting Uncertainty
Figure 7 illustrates how well different methods generalize a demonstrated trajectory across various surfaces. The methods compared include two trajectory reshaping techniques and two transportation methods, GPT and LWT, selected for their low out-of-distribution (o.o.d.) distortion. The figure shows that while Laplacian Editing often leads to highly distorted generalized trajectories, GPT and LWT perform similarly in terms of accuracy and computation time. For highly distorted surfaces (last column), some transported labels for the GP have a Jacobian determinant below zero, as evidenced by overlapping sections in the transported trajectory.
IV-B Multiple Reference Frames
Figure 8: Qualitative comparison of multi-reference frame parameterization. Comparison between HMM, TP-GMM, LEs, and our proposed method. The first row compares the performance in reproducing the training set demonstration, depicted as a dashed black line, where both HMM and TP-GMM are trained using all nine demonstrations, while LE and GPT are only trained with one demonstration (the central one) and generalized for each of the frames. In the second row, a random perturbation is applied to each frame, and the model is queried on the most likely trajectory. For GPT, the uncertainty in the generalization is depicted with the orange areas. The blue stars in LE and GPT are the points tracked during the motion, given that our proposed method does not rely on reference frames but only on source and target points. The yellow frame belongs to the source distribution.
In the literature, one of the main applications of task parameterization is the generalization w.r.t. one or more reference frames. For example, if we teach a robot how to pour water into a glass, we want the robot to automatically generalize the motion w.r.t. any glass position. The task, in this particular case, can be parameterized with the location of the reference frames of each object, which is necessary to track for a successful generalization of the motion. Typically, the motion is projected in any of the reference frames, and a policy is learned w.r.t. each of the frames, leaving out the decision on the relevance of each frame for every timestep. Task-Parametrized Gaussian Mixture Model (TP-GMM) learns a GMM model for the projected demonstration for each of the frames and, during executions, the Gaussians of each frame are combined using the property that the product of Gaussian (PoG) is still a Gaussian, see [2] for more details. Given the GMM, different control formulations are possible, for example only relying on the current state of the system, i.e., $\bm{\Delta}\bm{x}_{i}=\bm{f}(\bm{x}_{i})$ [53] [2] or by using a Hidden Markov Model (HMM) formulation that also considers the progress during the execution of the trajectory $\bm{x}_{i+1}=\bm{f}(\bm{x}_{i},\bm{\alpha}_{i})$, where $\bm{\alpha_{i}}$, selects the mean and variance that can be used in a tracking algorithm, such as a Linear Quadratic Regulator (LQR) [2, 54]. However, the latent transition matrix between the different states of the HMM is unknown. They need to be estimated using a forward pass algorithm, i.e., the Viterbi algorithm [54], that requires an initial guess trajectory to infer the most likely state transition that generated that initial guess and again generate the most likely motion according to the model. However, having an initial guess can be prohibitive when evaluating the movement in a novel configuration of starting and goal frames.
Unlike these task-parameterized approaches, our proposed method does not track only the reference frame but also a set of points that are relevant to the starting and goal objects. To describe a frame in 2D, we need at least two nonoverlapping points per frame; however, the transportation method will also scale if we extract more points; hence in Fig. 8, we extracted 5 points are tracked w.r.t. each reference frame when using GPT. However, for the LE, we used only 2 points per frame, as the redundancy of extracted points led to ambiguity in the assignment process and decreased algorithm performance. One of the main perks of our proposed method is the ability to generalize any motion skill generated by even only one demonstration, unlike GMM-based methods, where, to capture a meaningful mixture model, at least two diverse demonstrations need to be provided. Additionally, the GMM uncertainty of the final multi-frame model that results from the PoG does capture the uncertainty of the transportation, while, as depicted in Fig. 8, the GP transportation results in growing uncertainties when transporting points of the demonstrations that are less correlated with the source-target points. The illustrated trajectories (and uncertainties) both for GPT and LE are the result of the transportation of the single original demonstration.
Figure 9: Box plot results and performance ranking on frame configuration from the training set. The number on top of each box plot is the position of the method in the performance ranking. A higher value in the ranking (where the maximum is one) implies that the method has a more statistically significant discrepancy with another method than the next method in the ranking.
Fig. 8 highlights the discrepancy in the performance of GMM methods on the training set and the test set. At the same time, the reproduction of a known combination of the frames results in accurate rollouts of the policies both when executing them as a dynamical system (TP-GMM) [^1] Implementation available at https://github.com/BatyaGG/Task-Parameterized-Gaussian-Mixture-Model [53] or as an optimal tracking problem of a multi-transition Hidden Markov Model (HMM) [^2] Implementation available at https://gitlab.idiap.ch/rli/pbdlib-python/-/blob/master/notebooks/ [54]. However, when evaluating the test set generated on the random reorganization of the frames, the resulting trajectories do not successfully reach the goal frame either in position or orientation. To quantify and compare the different methods, we conducted a quantitative analysis comparing the generalization on known trajectories from the demonstration set or the reaching performance on a randomly generated frame set.
Quantitative Analysis
[…]
@@ IV-B Multiple Reference Frames
Considering that we have many models that can behave differently according to the amount and quality of the demonstration, it is not straightforward to deduce any conclusions on which method is statistically better from the boxplot of Figs. 9 and 10. For this reason, we run a U test, also known as the Mann-Whitney non-parametric test [56], to deduce if the distribution of results of each of the methods is statistically lower ($p<0.05$) than each of the others. When computing the U test between two methods, in case of a statistical difference, the winning method gets one point. The numbers on top of the figure for each of the methods indicate the performance ranking, i.e., the method that obtained the most points when computing the U test is going to be first in the ranking. When more methods share the same position in the ranking, it simply means they were significantly better than the same number of other methods during the comparisons.
@@
V Real Robot Validation
To validate the proposed method on real manipulation tasks, we selected three challenging tasks, i.e., robot reshelving (Sec. V-A), dressing (Sec. V-B) and cleaning (Sec. 15), to teach as a single demonstration and generalize it in different scenarios. These are all tasks where the training set will never be similar to the test set; for example, when dressing a human, the configuration and shape of the arm may change, and we expect the robot to generalize the behavior accordingly.
We controlled a Franka Robot using a Cartesian impedance control [^3] https://github.com/franzesegiovanni/franka_human_friendly_controllers. The motion policy is learned with a non-parametric function approximation for motor learning that uses as input the position and (a belief of) time proposed in [12]. The desired attractor position, orientation, stiffness, damping, and time belief update are learned as a function of the current position-time input. Our goal is to show how the proposed transportation theory can correctly generalize the pose, velocity, and stiffness of the robot. The following sections will summarize the robot validation experiments. A video of all the experiments can be found at https://youtu.be/bE6uOnAQBLo.
@@ V Real Robot Validation
V-A Robot Reshelving
Robot reshelving refers to picking an object in one location, moving it, and placing it in a desired position on a shelf. Our assumptions for the problem are:
  • one global frame movement primitive is learned from a single demonstration and transported in the different object/goal configurations;
  • corner points of the objects and the shelf slot are tracked rather than position/orientation. To describe the 3D pose of an object, we need at least 2 non-parallel vectors; hence, we need at least 3 non-overlapping and non-collinear points.
Fig. 11 depicts the experimental setup where a milk box, with an AprilTag [57] on it, has to be positioned on a compartment on a shelf, also marked by another tag. The tags are a handy way in robotics setups to track key points of objects. However, the source of the key point can be obtained with different vision models like [28, 58] without loss of generality for our proposed method. Before the demonstrations or execution, the robot searches for any tag in the spaces using the camera attached to its end-effector. For every tag, the transportation policy extracts a cube’s center and corners with predefined side dimensions as the markers. Fig. 11 shows how the demonstration for reshelving on the left of the central compartment can be generalized to any other floor, both on the left and right. We randomized the object position and orientation and the goal on the shelf ten different times, all successfully generalized. Table II shows the range of x,y,z, and yaw angles of the object and goal markers during the ten different executions, while Fig. 12 depicts the relative position w.r.t. the object and the frame of the different rollouts; from the figure it is possible to appreciate how the execution lines converge on the (initial) object position when picking and on the goal position when placing the object.
@@ V-A Robot Reshelving
Figure 12: Relative position of the end-effector w.r.t. the initial object and the goal position during multiple generalization rollouts in robot reshelving.
Figure 13: Dressing policy generalization. Cyan marker is the shoulder, magenta is the elbow, yellow is the wrist, and blue is the hand. The blue rollout end-effector trajectory starts from the red dot and finishes with the green dot. $\alpha$ is the angle of the elbow; the smaller the angle, the more complicated the generalization would be.
V-B Robot Dressing
@@
V-C Robot Surface Cleaning
@@
The point cloud is obtained from a RealSense D435 camera. The user selects the relevant area, after which the points are smoothed along the z-axis and organized into a grid on x and y coordinates of dimension $\times 20$ 20 where the corners are specified by a user input. This ensures that the source and target distributions have the same number of points, allowing for accurate matching and minimal estimation noise. One of the main advantages of the proposed method is that it does not need to reconstruct the surface but only learns the map from the source to the target point cloud. The transportation between the (smoothened) source and the (smoothened) target surface point cloud is modeled using a Stochastic Variational Gaussian Process Transportation (SV-GPT) to generalize the demonstrated policy position, orientation, and stiffness profile for a successful cleaning task. Given the large number of points in the source and target point cloud, i.e., 400 points, using a reduced set of inducing points, i.e., 100, makes fitting the transportation model more computationally efficient, see Appendix  A.
@@
Fig. 14 depicts the teaching of a cleaning task on a flat surface and the generalization on different heights, tilted, and curved surfaces that belong to common objects. The lower row shows what the robot perceives of the environment; the blue dots in space are the source distribution, recorded before giving the demonstration (depicted as dashed line), and target distributions recorded before executing the rollout transported policy (depicted a solid line). Fig. 14 also highlights how the rollouts follow the shape of the surface, showing a successful generalization of the robot position and orientation. As previously stated, no external force-torque sensor is used to adapt the orientation of the end-effector to the tangential direction of the surface. However, an observer of the applied external force between the robot and the surface is estimated from measured torques in the joints. Fig. 15 depicts the estimated norm of the force exchanged with the surface, where the same increasing/decreasing trend is captured on the different surfaces.
Figure 16: Food Stacking Task. The single demo is displayed on the left, and the motion generalization is on the right. The source and the target keypoints are automatically detected using DINO features correspondences [59]. The 3D location of the keypoint correspondence is obtained using the stereoscopic depth map.
V-D Towards policy transportation using DINO features
Relying on key points extracted from fiducial markers is not always practical in real-world applications. Nevertheless, given the raw camera input and an estimated depth, we can leverage new foundation models to automatically extract and find the corresponding features [59]. This experiment is similar to what was shown in DINObot [28]. However, we are not only roto-translating the policy, but we also apply the non-linear transformation proposed in this paper. Fig. 16 shows the source, recorded before giving the demonstration, and the target image on a new out-of-distribution configuration of the objects. A stereoscopic depth map is used to estimate the 3D location of each feature, so as to generate the necessary keypoints to perform the generalization. We ask the model to find a maximum of 7 key points, and a threshold of 0.2 was applied to the saliency maps, retaining only features with higher attention values; this avoids finding features on the table. We moved the objects around the workspace, always with the (toy) cheese on the left and the (toy) lettuce on the right, and used a transportation policy with a GP to perform the generalization of the original position and orientation labels. Despite very successful results in generalizing in 10 different configurations, we identify poor orientation generalization when swapping the location of the objects, i.e., lettuce on the left and cheese on the right side of the table. On the computational side, the main bottleneck is the DINO feature extraction, which, on the current implementation [59] on an RTX4060, can take almost a minute to find the keypoint correspondence. Moreover, the depth estimation turned out to be the brittle component of the pipeline when working with objects that are not flat.
VI Limitations and Open Challenges
@@
For example, we assume knowing the matching between the points of the source and the target distribution. However, in many complex scenarios, this limitation can be problematic, and some different pre-processing algorithms, such as optimal transport[60] or iterative closest point (ICP), or Coherent Point Drift [42] need to be adopted to perform the matching. Additionally, semantic matching can increase cross-domain generalization, for example, by adapting the reshelving strategy to a completely different shelf type or adjusting the dressing policy from an adult to a baby arm or to manipulate different food shapes or reshelving boxes of different dimensions.
Moreover, in the current implementation and experimentation, we assume full observability of the keypoints and can estimate their location with high accuracy. We foresee exciting directions in dealing with partial observability and different sensor noise, in particular when working with objects with very irregular shapes and estimating the keypoint location using the depth cameras.
@@
Another assumption of the developed method is that it deals with static environments, i.e., the target distribution does not change during the policy’s rollout. However, this assumption can fail when dealing with the reshelving of moving objects or when trying to dress real humans that will probably move before and during the interaction [61, 62]. Nevertheless, supposing to know the state of the target distribution, the transportation policy can be updated inexpensively online since only the label $\hat{y}$ of Eq. (2) and the (cheap) linear transformation needs to be recomputed. However, the fitting of the transported policy $\hat{f}$ makes it challenging to perform the generalization online.
@@
Finally, given that in complex scenarios, the generalization may be inaccurate, the use of interactive human corrections may increase the resulting manipulation performance [1]. However, changing the generalized policy opens the question of whether interactive corrections should be propagated back to the original policy labels and how. Additionally, in case many source distributions/policies are recorded, selecting the best one, according to some similarity to the current observed target distribution, can open exciting developments of scaling the proposed theory to work on a big database of keypoint parameterized demonstrations.
VII Conclusions
@@
In this paper, we address the prominent but challenging problem of policy generalization to novel unseen task scenarios. We formulate a novel policy transportation theory that, given a set of matched source and target points in the task space of the robot, regresses the function that, most likely, would match the source and target distribution. Additionally, we showed how the same transportation function and its derivatives can be exploited to transport the original policy dynamics, rotation, and stiffness while retaining uncertainties in the process.
The same transportation algorithm was tested with different state-of-the-art regressors and compared with different generalization methods, showing how, even with only one demonstration, it results in better or comparable motion generalization.
@@
We validated the proposed approach on a Franka Robot, testing it on four different tasks: product reshelving, arm dressing, and surface cleaning, and table-top pick-and-place. These various tasks were never tackled together by the same generalization algorithm, and they were usually performed with ad hoc solutions, for example, to keep a constant force when cleaning a surface. Despite this, our policy transportation algorithm performed successfully in all of them. The tracking requirements were satisfied using fiducial markers or directly the point cloud estimated with a stereo camera. Future development will have to focus on scaling the manipulation generalization using large, unmatched, and partially observable point clouds of complex, deformable, and moving objects while allowing the use of human feedback in the fine-tuning of the resulting policy, asking for human attention when facing high-uncertainty generalization regions, or actively searching for missing keypoints that would reduce the total transportation uncertainty.
Appendix A Gaussian Process Regression
where $\bm{X}_{*}\in\mathbb{R}^{n\times m}$ are the evaluation inputs and $\bm{X}\in\mathbb{R}^{N\times m}$ and $\bm{y}\in\mathbb{R}^{N\times l}$ are the training inputs and outputs; the correlations $\bm{K}_{\bm{X},\bm{X}}\in\mathbb{R}^{N\times N}$, $\bm{K}_{\bm{X}_{*},\bm{X}}\in\mathbb{R}^{n\times N}$, $\bm{K}_{\bm{X}_{*},\bm{X}_{*}}\in\mathbb{R}^{n\times n}$ are the correlation matrix between every input training point, between testing and training points and between testing points. Every entry of this matrices is computed using a kernel function, for example, the squared exponential kerneldefined as
where $\bm{x_{i}},\bm{x_{j}}$ are any pair of training-training, testing-training, or testing-testing data points; $\sigma_{p}$ and $\ell$ are the kernel hyperparameters and, together with the likelihood noise $\sigma_{n}$ are optimized with a maximization of the log-likelihood of the given data [48]. In the context of this paper, when dealing with multi-output GPs, we assume the same smoothness, prior noise and likelihood noise among the models, hence we compute $(\bm{K}_{\bm{X},\bm{X}}+\sigma^{2}_{n}\bm{I})^{-1}$ once for all the three models and during inference the vector $\bm{K}_{\bm{X}_{*},\bm{X}}$ is also the same for the three outputs.
Moreover, given the computationally expensive inversion of the covariance matrix that grows with the number of data points, optimizing kernel hyperparameters and inference becomes prohibitive when having big datasets. Therefore, variational inference aims to approximate the true posterior $p(\bm{f}|\bm{X},\bm{y})$ with a simpler distribution $q(\bm{f})$. Stochastic Variational Gaussian Process (SVGP) optimizes the location and values of a set of pseudo data, usually referred to as inducing points $\bm{Z}$ and a variational distribution $q(\bm{u})$. The parameters of the kernel and the variational parameter of the approximated distribution are optimized using the evidence lower bound (ELBO) as the evidence of the data[63, 64] .
The use of SVGP, in the context of this paper, can make the computation significantly faster when fitting a transportation function with many points from the source and target distribution; however, any of the proposed algorithms, formalized in this paper is independent of the approximation nature of the GP.
When dealing with derivatives of transportation maps that are estimated with probabilistic functions, it is necessary to estimate the uncertainty over the predicted derivatives. Fortunately, a GP derivative is also a GP [48] and its existence will depend on the differentiability of the kernel function. The correlation between derivative samples can be expressed as the second partial derivative $k^{11}=\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}k(x_{i},x_{j})$ while the correlation between derivative samples and function samples is $k^{10}=\frac{\partial}{\partial x_{i}}k(x_{i},x_{j})$. Thus, the mean and variance prediction of the derivative of the Gaussian Process become
[…]
[Uncaptioned image] Giovanni Franzese obtained a Ph.D. in Robotics from the Department of Cognitive Robotics at TU Delft in 2025. He received a BSc degree (2016) in Mechanical Engineering and an MSc degree (2018) in Mechatronics and Robotics at Politecnico di Milano, Italy. He is an ELLIS member for artificial intelligence. His research focuses on Interactive Imitation Learning applied to robot manipulation.
[Uncaptioned image] Ravi Prakash is an Assistant Professor at the department of Cyber Physical Systems (CPS), IISc Bangalore, India. Prior to this, he was a Post-Doctoral researcher in the Learning and Autonomous Control group in the department of Cognitive Robotics, TU Delft. He earned his Ph.D.(2022) in Control and Automation at the Department of Electrical Engineering, Indian Institute of Technology Kanpur. He is a DAAD AInet Postdoc fellow for AI and Robotics.
[Uncaptioned image] Cosimo Della Santina is currently an Associate Professor with TU Delft, Delft, The Netherlands, and a Research Scientist with the German Aerospace Institute (DLR). He received the Ph.D. degree (cum laude) in robotics from the University of Pisa, Italy, in 2019. From 2017 to 2019, he was a visiting Ph.D. student and a Postdoc at the CSAIL at the Massachusetts Institute of Technology. His research interest includes providing motor intelligence to physical systems, focusing on elastic and soft robots. He received the Georges Giralt Ph.D. Award in 2020, the IEEE RAS Early Academic Career Award in 2023, an ERC StG, and an NWO VENI.
[Uncaptioned image] Jens Kober is an associate professor at the TU Delft, Netherlands. He worked as a postdoctoral scholar jointly at the CoR-Lab, Bielefeld University, Germany and at the Honda Research Institute Europe, Germany. He graduated in 2012 with a PhD Degree in Engineering from TU Darmstadt and the MPI for Intelligent Systems. For his research he received the annually awarded Georges Giralt PhD Award for the best PhD thesis in robotics in Europe, the 2018 IEEE RAS Early Academic Career Award, the 2022 RSS Early Career Award, and has received an ERC Starting grant.